Peter's age 13 years ago was 1/3 of his age seven years hence. How old is peter?
Answer by stanbon(75887)

Peter's age 13 years ago was 1/3 of his age seven years hence. How old is peter?
------
Equation::
p-13 = (1/3)(p+7)
----
3p - 39 = p + 7
2p = 46
Peter's age = 23 yrs.
=======================
Cheers,
Stan H.